### v1

@@ Abstract
In this work, we investigate the discrete Calderón problem on grid graphs of dimension three or higher, formed by hypercubic structures. The discrete Calderón problem is concerned with determining whether the discrete Dirichlet-to-Neumann (DtN) operator, which links boundary potentials to boundary current responses, can uniquely identify the conductivity values on the graph edges. We provide an affirmative answer to the question, thereby extending the classical uniqueness result of Curtis and Morrow for two-dimensional square lattices. The proof employs a novel slicing technique that decomposes the problem into lower-dimensional components. Additionally, we support the theoretical finding with numerical experiments that illustrate the effectiveness of the approach.
@@ 1 Introduction
The Calderón problem named after Alberto Calderón falls into the class of inverse boundary value problems, mostly investigated in an open bounded domain $\Omega\subseteq\mathbb{R}^{d}$, and focuses on the recovery of the interior conductivity only by the measurements on the boundary $\partial\Omega$. Physically, the interior electric potential follows Kirchhoff’s current law, which takes the form of a second-order elliptic partial differential equation (PDE) in the divergence form. Since the seminal work of Alberto Calderón [12], the problem has been extensively studied in the last fourty years, with many deep theoretical results [5, 27].
The “discrete” Calderón problem focuses on a finite graph $G=(E,D,\partial D)$. The set $E$ contains edges that connect two nodes in the set $D\cup\partial D$. The electric system on the graph can be viewed as connecting homogeneous electric wires according to the graph structure $G$. Then the electric potential function, on both the edges and nodes, that is, the wires and their connecting points, is fully determined by its value on nodes $D\cup\partial D$, and by the linear interpolation of nodal values for each edge.
These two versions of the Calderón problem (continuous versus discrete) represent two fundamentally different ways to formalize the physical problem. In a series of works, Borcea and her collaborators [6, 7, 8, 9, 10] established deep connections between the continuous Calderó problem with the discrete one. However, the mathematical tools utilized to investigate these two versions differ significantly. In particular, very different excitations (i.e., the boundary potential chosen to induce the current measurement) are employed in the analysis. For the continuous Calderón problem, the so-called complex geometric optics solutions [26] with high-frequency oscillations represent one powerful tool to derive theoretical results. In contrast, in the discrete case, the high-frequency property is inherently incompatible with the finite graph structure, and a new type of excitation has been employed, i.e., excitations that are able to localize all interior potential within a sub-domain of the graph [19].
In a graph $G$, $D$ and $\partial D$ are finite sets with no common elements, and $E:=\{pq:p,q\in D\cup\partial D\}$ where $pq$ is an unordered pair. For any node $p\in D\cup\partial D$, its neighbouring nodes $\mathcal{N}(p)$ is defined by the set $\{q:pq\in E\}$. Each edge $pq$ in the graph can also be assigned a positive conductivity value $\gamma$, where by definition it holds that $\gamma_{pq}=\gamma_{qp}>0$. Moreover, real-valued functions on a finite set $U$ can be identified with vectors in $\mathbb{R}^{U}$, in which way we define the potential function (or vector) as $\mathbf{u}\in\mathbb{R}^{D\cup\partial D}$.
Physically, the interior potential $\mathbf{u}$ satisfies Kirchhoff’s current law, and mathematically, we express it as the discrete Laplacian operator $\Delta_{\gamma}$, defined by
@@
where $\varphi\in\mathbb{R}^{\partial D}$ is the Dirichlet boundary condition. The unique solvability of (1.1) is well known [19] for general graphs. Then one can define an injective solution operator $S_{\gamma}$ by
@@
where $\mathbf{u}$ is the unique solution of (1.1) with the Dirichlet condition $\varphi$. Assume that each boundary vertex $p\in\partial D$ is connected to precisely one interior vertex, that is, $\mathcal{N}(p)=\{q\}\subseteq D$. Then the boundary current data can be represented as
@@
The DtN matrix $\Lambda_{\gamma}$ is then the composition of the two operators:
@@
Then the discrete Calderón problem reads: Given the DtN matrix $\Lambda_{\gamma}$, is it possible to uniquely recover the conductivity $\gamma$? In the literature it is also known as the discrete inverse conductivity problem. In their seminal work [19], Curtis and Morrow established the uniqueness of the solution for the inverse problem in two-dimensional square lattices. This study aims to extend the result of Curtis and Morrow [19] to three and higher dimensional lattice structures.
Now we precisely define the graph $G=(E,D,\partial D)$ in a high-dimensional grid. The set of interior vertices $D$ consists of all integer tuples $(x_{i})_{i=1}^{d}\in\mathbb{Z}^{d}$ such that each coordinate adheres to the constraints $1\leq x_{i}\leq n$. This can be expressed as
@@
The boundary vertex set $\partial D$ is comprised of nodes for which exactly one coordinate is equal to either 0 or $n+1$, while the remaining coordinates range over the set $\{1,\ldots,n\}$:
$\partial D$ $\coloneqq\{p\in\mathbb{Z}^{d}:d(p,D)=\min_{q\in D}\|q-p\|_{\ell^{1}}=1\}.$ (1.5)
The edge set $E$ consists of unordered pairs $\{p,q\}$ where $p=(p_{i})_{i=1}^{d}$ and $q=(q_{i})_{i=1}^{d}$ are vertices in $D\cup\partial D$ that meet the following criteria: (i) $\|p-q\|_{\ell^{1}}=1$; and (ii) The pair does not reside entirely within the boundary, i.e., $\{p,q\}\not\subseteq\partial D$. This defines the multi-dimensional hypercubic lattice graphs:
@@
The main contribution of this work is the following uniqueness result for the discrete Calderón problem on multi-dimensional hypercubic graphs.
@@
For the multi-dimensional grids graph $G$ given in (1.6), the DtN matrix $\Lambda_{\gamma}$ uniquely determines the conductivity $\gamma$ on each edge $E$.
@@
Theorem 1.1 generalizes the result of Curtis and Morrow [19] on the uniqueness of the discrete Calderón problem for two-dimensional square lattice to the multi-dimensional case. The proof of Theorem 1.1 employs mathematical induction and a slicing technique that effectively breaks down the complex lattice into more manageable components. The uniqueness of the conductivity $\gamma$ is established through an inductive procedure that is applied slice by slice, based on the slice decomposition strategy. The inductive step based on refined characterization of the algebraic properties of the solution spaces $\mathcal{U}^{(t)}$ on hypercubic lattices. Moreover, the analysis strategy is constructive, and lends itself to an efficient algebraic reconstruction algorithm that can be used to recover the conductivity $\gamma$ from the DtN matrix $\Lambda_{\gamma}$. In Section 5, we present the reconstruction algorithm and numerical illustrations to show the feasibility of the reconstruction.
@@
Last we discuss related works on the discrete Calderón problem. The problem in the two-dimensional case has been extensively studied since the seminal work of Curtis and Morrow [19]. There are many interesting theoretical results for the discrete Calderón problem on planar graphs, including square lattices [19, 20] and circular planar graphs [17, 15, 18, 22]. The work [14] proved a uniqueness result from one boundary measurement under a certain monotonicity assumption. Boyer et al [11] investigated the recovery of a complex-valued conductivity (or complex-valued potential) from the DtN matrix $\Lambda_{\gamma}$, and established the unique recovery up to a set of measure zero.
@@
In sharp contrast, the discrete Calderón problem in the multi-dimensional case has only been scarcely studied so far. Lam and Pylyavskyy [23] investigated the recovery of the conductivity from the so-called $R$-response matrix, proved that the problem generally does not have uniqueness, and established uniqueness for a certain class of purely cylindrical networks. The analysis in [23] relies heavily on grove combinatorics. Araúz et al [1] treated cubic lattices for a particular excitation, and proved the unique recovery of the conductivity from the Robin-to-Neumann map. The proof relies on proving that certain overdetermined boundary value problems on graphs admit a unique solution. Compared with the work [1], we employ the space of admissible excitations as the main tool, and the approach extends readily to general multi-dimensional hypercubic lattice graphs and possibly more general shaped graphs. Moreover, the proposed approach is constructive in the sense that it lends itself directly to an algebraic reconstruction algorithm similar to Curtis-Morrow algorithm for square lattices [20].
@@
The works [24, 13, 21, 25] investigated a closely related inverse problem of recovering the potential in the discrete Schrödinger operator from the (partial) DtN type matrix. Note that due to the lack of a neat Liouville type transformation on graphs, the discrete inverse Schrödinger problem is not equivalent to the discrete Calderón problem. Interestingly, in the work [21], the slicing decomposition was also crucially employed in the uniqueness proof. Although this work shares the spirit with [21], the recovery of the conductivity $\gamma$ requires several additional investigations of the algebraic properties of the solution space $\mathcal{U}^{(t)}$, and thus the analysis is much more intricate. Also we refer interested readers to [4, 3, 16] for other recent important works on inverse problems on graphs.
@@
The rest of the paper is organized as follows. In Section 2, we construct a specifically designed boundary potential that confines nonzero currents to a corner region. In Section 3, we collect three fundamental topological properties of grid graphs, which are essential for the proof of Theorem 1.1. In Section 4, we present the proof of Theorem 1.1. The proof leads to an explicit reconstruction algorithm that incrementally recovers the conductivity $\gamma$, slice by slice, relying solely on the DtN matrix $\Lambda_{\gamma}$. Finally, in Section 5, we present numerical results to illustrate the algorithm derived from the uniqueness proof. Throughout this work, we denote the space of real-valued functions defined on a set $U$ by $\mathbb{R}^{U}$. We extend functions defined on subsets $U\subset V$ to functions on $V$ by assuming they are zero outside $U$ and then we can define the inclusion $\mathbb{R}^{U}\subseteq\mathbb{R}^{V}$. Given a matrix $A:\mathbb{R}^{V_{1}}\rightarrow\mathbb{R}^{V_{2}}$, we define its sub-matrix with column indices $U_{1}\subseteq V_{1}$ and row indices $U_{2}\subseteq V_{2}$ by $A(U_{2};U_{1})$.
@@ 2 Corner excitations
In this section, we introduce a special type of boundary excitations that induce well localized potentials and give their important properties. This is the main technical tool to carry out with the inductive steps along the slices. In the proof, the key step is inductive, starting from a subdomain with the known conductivity, to recover the conductivity on a larger region. That is, the domain with known conductivity grows with the induction step, which eventually covers the entire edge set. To rigorously formulate the inductive step, we define a class of growing regions slice by slice as
@@
The inductive step is that, given the conductivity $\gamma$ on $E(L_{t}^{\mathcal{S}}\cup K_{t}^{\mathcal{S}},L_{t}^{\mathcal{S}}\cup K_{t}^{\mathcal{S}})$, recover $\gamma$ on $E(L_{t+1}^{\mathcal{S}}\cup K_{t+1}^{\mathcal{S}},L_{t+1}^{\mathcal{S}}\cup K_{t+1}^{\mathcal{S}})$. The analysis requires also considering a subset $J_{t}^{\mathcal{S}}$ on the boundary that lies in the lower part of the graph $G$ divided by $L_{t+1}$:
$K_{t}^{+}$ $=\{(x_{i})_{i}\in K_{t}:\max_{i}x_{i}=n+1\},$ $K_{t}^{\mathcal{S}+}$ $=\bigcup_{\ell=0}^{t}K_{\ell}^{+},$ (2.2a)
$K_{t}^{-}$ $=\{(x_{i})_{i}\in K_{t}:\min_{i}x_{i}=0\},$ $K_{t}^{\mathcal{S}-}$ $=\bigcup_{\ell=0}^{t}K_{\ell}^{-}.$ (2.2b)
@@
We refer to Fig. 1 for a schematic illustration of the interface partition in the three-dimensional case.
@@
Intuitively, to recover the conductivity on the current sub-region $L_{t+1}^{\mathcal{S}}\cup K_{t+1}^{\mathcal{S}}$, we select a set of boundary excitations that ”focus” on the sub-region and use its boundary measurements to finish the inductive step. That is, the non-zero current is localized in the sub-region.
@@
The excitations are exactly in the kernel of the following map
We can factorize the map $T^{(t)}$ into two parts: one is the map from the potential supported at the boundary $J_{t}^{\cal S}$ to the potential supported on the boundary $L_{t+1}$, and the other is the map from the potential supported on $L_{t+1}$ to the boundary currents at $\partial D\setminus J_{t}^{\cal S}$. We refer readers to Fig. 2 for a schematic illustration in the three-dimensional case: the sought-for excitations are precisely supported on the blues nodes with the boundary current vanishing at yellows nodes. More precisely, the map $T_{t}^{(1)}$ is defined by
@@
For $T_{t}^{(2)}$, we first introduce the upper subgraph subgraph $G^{(t)}$ above the interface $L_{t+1}$:
$G^{(t)}=\bigl(E^{(t)},\,D\setminus L_{t+1}^{\mathcal{S}},\,(L_{t+1}\cup\partial D)\setminus J_{t}^{\mathcal{S}}\bigr),$ (2.5)
where the set $E^{(t)}$ contains all edges in the set $E$ connecting nodes within $D\setminus L_{t}^{\mathcal{S}}$. The unique solvability of discrete Dirichlet problem on graph $G^{(t)}$ holds, in view of [19, Remark 3] and hence the DtN map $\Lambda^{(t)}$ on subgraph $G^{(t)}$ is well defined. We can now describe the map $T_{2}^{(t)}$, from the potential at $L_{t+1}$ to the current at $\partial D\setminus J_{t}^{\cal S}$, by a submatrix of $\Lambda^{(t)}$:
$T_{2}^{(t)}:=\Lambda^{(t)}(\partial D\setminus J_{t}^{\mathcal{S}};L_{t+1}):\mathbb{R}^{L_{t+1}}\to\mathbb{R}^{\partial D\setminus J_{t}^{\mathcal{S}}}.$ (2.6)
The next result rigorously establish the factorization.
@@
For any $t$ in the range $d-1\leq t\leq dn-1$, the operator $T^{(t)}$ can be factorized as $T^{(t)}=T_{2}^{(t)}\circ T_{1}^{(t)}.$ (2.7)
@@
Fix any arbitrary $\varphi\in\mathbb{R}^{J_{t}^{\mathcal{S}}}$, and consider the zero extension of $\varphi$ to $\partial D$, still denoted by $\varphi$. The corresponding interior potential is given by $\mathbf{u}:=S_{\gamma}\varphi$. By its definition, $\mathbf{u}$ is $\gamma$-harmonic on $G$, and also within the subgraph $G^{(t)}$. By the definition of the operator $T_{1}^{(t)}$, we have
@@
Since the discrete Dirichlet problem on $G^{(t)}$ has a unique solution, the restriction of $\mathbf{u}$ to $G^{(t)}$ can be represented by
$\mathbf{u}=S_{\gamma}^{(t)}\bigl(T_{1}^{(t)}\varphi\bigr)\quad\text{in }G^{(t)},$
where $S_{\gamma}^{(t)}$ denotes the solution operator on $G^{(t)}$. Now we compute the boundary current and arrive at
@@
The factorization in Proposition 2.1 leads directly to the kernel inclusion:
@@
The identity $\ker T_{1}^{(t)}=\ker T^{(t)}$ then follows from injectivity of $T_{2}^{(t)}$, which will be proved in Section 3.
In fact, at the slice $t$, we impose boundary potentials from the kernel of $T_{1}^{(t)}$, record the corresponding boundary currents, and then recover the conductivity between the layers $L_{t}\cup K_{t}$ and $L_{t+1}\cup K_{t+1}$. Since the boundary potentials are chosen from $\ker T_{1}^{(t)}$, the interior potentials are localized within the set $J_{t}^{\mathcal{S}}\cup L_{t}^{\mathcal{S}}$, by the definition of the mapping $T_{1}^{(t)}$. We now define
@@
where $S_{\gamma}$ is the solution operator defined in (1.2), and the notation $\big|_{J_{t}^{\mathcal{S}}\cup L_{t}^{\mathcal{S}}}$ denotes the restriction of functions from the domain $D$ to the region $J_{t}^{\mathcal{S}}\cup L_{t}^{\mathcal{S}}$, on which the elements $\mathbf{u}\in\mathcal{U}^{(t)}$ are naturally supported. Note that, when recovering the conductivity between $L_{t}\cup K_{t}$ and $L_{t+1}\cup K_{t+1}$, only the boundary potentials in the quotient space
@@
are relevant for the reconstruction subproblem at step $t$. The quotient makes sense in view of the following result of inclusion in the sense $\mathbb{R}^{J_{t-1}^{\mathcal{S}}}\subseteq\mathbb{R}^{J_{t}^{\mathcal{S}}}$.
@@
For $\varphi\in\ker T_{1}^{(t-1)}$, let $\mathbf{u}:=S_{\gamma}\varphi$. By construction,
$\mathbf{u}=0\quad\text{in }D\setminus L_{t}\supseteq L_{t+1},$
@@
The monotonicity of $\mathcal{U}^{(t)}$ follows directly:

### v2

@@ Abstract
In this work, we investigate the discrete Calderón problem on (hyper)cubic grid graphs in dimension three or higher. The discrete Calderón problem is concerned with determining whether the discrete Dirichlet-to-Neumann matrix, which links boundary potentials to boundary current responses, can uniquely determine the conductivity values on the graph edges. We provide an affirmative answer to the question, thereby extending the classical uniqueness result of Curtis and Morrow for two-dimensional square lattices. The proof employs a slicing technique that decomposes the problem into lower-dimensional components. Additionally, we support the theoretical finding with numerical experiments that illustrate the effectiveness of the approach.
@@ 1 Introduction
The Calderón problem named after Alberto Calderón falls into the class of inverse boundary value problems, mostly investigated in an open bounded domain $\Omega\subseteq\mathbb{R}^{d}$, and focuses on the recovery of the interior conductivity only by the measurements on the boundary $\partial\Omega$. Physically, the interior electric potential satisfies Kirchhoff’s current law, which takes the form of a second-order elliptic partial differential equation in the divergence form. Since the seminal work of Alberto Calderón [12], the Calderón problem has been extensively studied in the last forty years, with many deep theoretical results [10, 29, 21].
The discrete Calderón problem focuses on a finite graph $G=(E,D,\partial D)$. The set $E$ contains edges that connect two nodes in the set $D\cup\partial D$. Roughly speaking, the electric system on the graph $G$ can be viewed as connecting homogeneous electric wires according to the graph structure $G$. Then the electric potential on the edges and nodes, that is, the wires and their connecting points, is fully determined by its value on nodes $D\cup\partial D$, and by a linear interpolation of nodal values for each edge.
These two versions of the Calderón problem (continuous versus discrete) represent two fundamentally different ways to formalize the physical problem. The mathematical tools utilized to investigate these two versions differ significantly. In particular, very different excitations (i.e., the boundary potential chosen to induce the current measurement) are employed in the analysis. For the continuous Calderón problem, the so-called complex geometric optics solutions [28] with high-frequency oscillations represent one powerful tool to derive theoretical results. In contrast, in the discrete case, the high-frequency property is inherently incompatible with the finite graph structure, and a new type of excitation has been devised, i.e., excitations that are able to localize all interior potential within a sub-domain of the graph [19].
Let $G$ be a graph with vertex set $\overline{D}=D\cup\partial D$, where $D$ and $\partial D$ are disjoint finite sets. Its edge set is $E=\{pq:pq\in\overline{D}\}$, the unordered pairs on $\overline{D}$. For each vertex $p\in\overline{D}$, the neighbor set is $\mathcal{N}(p)=\{q\in\overline{D}:pq\in E\}$. Every edge $pq$ is assigned a positive conductivity $\gamma_{pq}=\gamma_{qp}>0$. Real-valued functions on any set $V$ are identified with vectors in $\mathbb{R}^{V}$, and we denote the potential function by $\mathbf{u}\in\mathbb{R}^{\overline{D}}$.
Physically, the potential $\mathbf{u}$ satisfies Kirchhoff’s current law on the interior nodes, and mathematically, we express it via the discrete Laplacian operator $\Delta_{\gamma}$, defined by
@@
where $\varphi\in\mathbb{R}^{\partial D}$ is the Dirichlet boundary condition. The unique solvability of problem (1.1) is well known for general graphs [19]. Then one can define an injective solution operator $S_{\gamma}$ by
@@
where $\mathbf{u}$ is the unique solution of problem (1.1) with the Dirichlet condition $\varphi$. Assume that each boundary vertex $p\in\partial D$ is connected to precisely one interior vertex, that is, $\mathcal{N}(p)=\{q\}\subseteq D$. Then the boundary current data can be represented as
@@
The Dirichlet-to-Neumann (DtN) matrix $\Lambda_{\gamma}:\mathbb{R}^{\partial D}\to\mathbb{R}$ is the composition of these two operators:
@@
Then the discrete Calderón problem reads: Given the DtN matrix $\Lambda_{\gamma}$, is it possible to uniquely identify the conductivity $\gamma$? In the literature it is also known as the discrete inverse conductivity problem. In their seminal work [19], Curtis and Morrow established the uniqueness of the discrete Calderón problem on two-dimensional square lattices. This study aims to extend the result of Curtis and Morrow [19] to three and higher dimensional lattice structures.
Now we precisely define the graph $G=(E,D,\partial D)$ in a high-dimensional grid. The set $D$ of interior nodes consists of all integer tuples $(x_{i})_{i=1}^{d}\in\mathbb{Z}^{d}$ such that each coordinate $x_{i}$ adheres to the constraints $1\leq x_{i}\leq n$. This can be expressed as
@@
The set $\partial D$ of boundary vertices is comprised of nodes for which exactly one coordinate is equal to either 0 or $n+1$, while the remaining coordinates range over the set $\{1,\ldots,n\}$:
$\partial D\coloneqq\{p\in\mathbb{Z}^{d}:d(p,D):=\min_{q\in D}\|q-p\|_{\ell^{1}}=1\},$ (1.5)
where $\|\cdot\|_{\ell^{1}}$ denotes the $\ell^{1}$ norm of vectors. The edge set $E$ consists of unordered pairs $\{p,q\}$, where $p=(p_{i})_{i=1}^{d}$ and $q=(q_{i})_{i=1}^{d}$ are vertices in $D\cup\partial D$ that meet the following two criteria: (i) $\|p-q\|_{\ell^{1}}=1$; and (ii) The pair does not reside entirely within the boundary, i.e., $\{p,q\}\not\subseteq\partial D$. This defines a multi-dimensional hypercubic lattice graph:
@@
The main result of this work is the following uniqueness result for the discrete Calderón problem on multi-dimensional hypercubic graphs.
@@
For the multi-dimensional grid graph $G$ given in (1.6), the DtN matrix $\Lambda_{\gamma}$ uniquely determines the conductivity $\gamma$.
@@
Theorem 1.1 extends the result of Curtis and Morrow [19] on the uniqueness of the discrete Calderón problem for two-dimensional square lattices to multi-dimensional hypercubic graphs. The proof of Theorem 1.1 employs mathematical induction and a slicing technique that effectively breaks down the complex lattice into more manageable components. The uniqueness of the conductivity $\gamma$ is established through an inductive procedure that is applied slice by slice, based on the slice decomposition strategy. The inductive step is based on refined characterizations of the algebraic properties of the solution spaces $\mathcal{U}^{(t)}$ on hypercubic lattices. Moreover, the analysis strategy is constructive, and lends itself to an efficient algebraic reconstruction algorithm that can be used to recover the conductivity $\gamma$ from the DtN matrix $\Lambda_{\gamma}$. In Section 5, we present the reconstruction algorithm and numerical illustrations to show the feasibility of the reconstruction.
@@
Finally, we discuss related works on the discrete Calderón problem. The problem in the two-dimensional case has been extensively studied since the seminal work of Curtis and Morrow [19]. There are many interesting theoretical results for the discrete Calderón problem on planar graphs, including square lattices [20] and circular planar graphs [18, 15, 17, 24]. The work [14] proved a uniqueness result from one boundary measurement under a certain monotonicity assumption. Boyer et al [11] investigated the recovery of a complex-valued conductivity (or complex-valued potential) from the DtN matrix $\Lambda_{\gamma}$, and established the unique recovery up to a set of measure zero. From the DtN matrix on a tree graph, Gernandt and Rohleder [22] proved an explicit formula which relates the DtN matrix to the pairwise weighted distances of the leaves of the tree (and thus the weighted tree).
@@
The discrete Calderón problem in the multi-dimensional / non-planar case has only been scarcely studied so far. On cylindrical networks, Lam and Pylyavskyy [25] investigated the recovery of the conductivity from the so-called $R$-response matrix, proved that the problem generally does not have uniqueness, and established uniqueness for a certain class of purely cylindrical networks. The analysis in [25] relies heavily on grove combinatorics. Araúz et al [1] treated cubic lattices for a particular excitation, and proved the unique recovery of the conductivity from the Robin-to-Neumann map. The proof relies on proving that certain overdetermined boundary value problems on graphs admit a unique solution. Compared with the work [1], we employ the space of admissible excitations as the main tool, and the approach extends readily to general multi-dimensional hypercubic lattice graphs and possibly graphs of more general shapes. Moreover, the proposed approach is constructive in the sense that it lends itself directly to an algebraic reconstruction algorithm similar to the Curtis-Morrow algorithm for square lattices [20].
@@
The works [26, 13, 23, 27] investigated a closely related inverse problem of recovering the potential in the discrete Schrödinger operator from the (partial) DtN type matrix. Note that due to the lack of a neat Liouville type transformation on graphs, the discrete inverse Schrödinger problem is not equivalent to the discrete Calderón problem. Interestingly, in the work [23], the slicing decomposition was also crucially employed in the uniqueness proof. Although this work shares the spirit with [23], the recovery of the conductivity $\gamma$ requires several additional investigations of the algebraic properties of the solution spaces $\mathcal{U}^{(t)}$, and thus the analysis is much more intricate. Also we refer interested readers to [4, 3, 16] for other recent contributions on inverse problems on graphs.
Formally, the discrete Calderón problem can be viewed as a discretization of the continuous version. By drawing on this connection and suitable discretization schemes, Borcea and her collaborators [8, 7, 6, 5, 9] have developed several effective numerical schemes for the continuous Calderón problem. Nonetheless, to the best of our knowledge, rigorously establishing the connection of the uniqueness and stability results between these two versions of the Calderón problem is still missing.
@@
The rest of the paper is organized as follows. In Section 2, we construct a specifically designed boundary potential that confines nonzero currents to a corner region. In Section 3, we collect three fundamental topological properties of grid graphs, which are essential for the proof of Theorem 1.1. In Section 4, we present the proof of Theorem 1.1. The proof leads to an explicit reconstruction algorithm that incrementally recovers the conductivity $\gamma$, slice by slice, relying solely on the DtN matrix $\Lambda_{\gamma}$. Finally, in Section 5, we present numerical results to illustrate the algorithm derived from the uniqueness proof.
Throughout this work, we denote the space of real-valued functions defined on a set $U$ by $\mathbb{R}^{U}$. We extend functions defined on subsets $U\subset V$ to functions on $V$ by assuming they are zero outside $U$ and then we can define the inclusion $\mathbb{R}^{U}\subseteq\mathbb{R}^{V}$. Given a matrix $A:\mathbb{R}^{V_{1}}\rightarrow\mathbb{R}^{V_{2}}$, we define its sub-matrix with column indices $U_{1}\subseteq V_{1}$ and row indices $U_{2}\subseteq V_{2}$ by $A(U_{2};U_{1})$. Given a subset $U\subset\overline{D}$, the notation $\big|_{U}$ denotes the restriction of functions from the domain $\overline{D}$ to the region $U$, and the convention also applies to vectors defined on the edge set $E$. For any two subsets $V_{1},V_{2}\subset\overline{D}$, we denote by $E(V_{1},V_{2})$ the subset of the set $E$ comprised of edges starting from the vertex set $V_{1}$ and ending at $V_{2}$.
@@ 2 Corner excitations
In this section, we describe a special type of boundary excitations that induce well localized potentials and give their important properties. This is the main technical tool to carry out the inductive steps along the slices. The inductive step starts from a subdomain with the known conductivity, and then aims to recover the conductivity on a larger region. That is, the domain with known conductivity grows with the induction step, which eventually covers the entire edge set $E$. To rigorously formulate the inductive step, we define a class of growing regions slice by slice as
@@
The inductive step is that, given the conductivity $\gamma$ on the edge set , we recover $\gamma$ on the enlarged edge set $E(L_{t}^{\mathcal{S}},\overline{D})$. The analysis requires also considering a subset $J_{t}^{\mathcal{S}}$ on the boundary $\partial D$ that lies in the lower part of the graph $G$ divided by $L_{t+1}$:
$K_{t}^{+}$ $=\{(x_{i})_{i=1}^{d}\in K_{t}:\max_{i}x_{i}=n+1\},$ $K_{t}^{\mathcal{S}+}$ $=\bigcup_{\ell=0}^{t}K_{\ell}^{+},$ (2.2a)
$K_{t}^{-}$ $=\{(x_{i})_{i=1}^{d}\in K_{t}:\min_{i}x_{i}=0\},$ $K_{t}^{\mathcal{S}-}$ $=\bigcup_{\ell=0}^{t}K_{\ell}^{-}.$ (2.2b)
@@
We refer the reader to Fig. 1 for a schematic illustration of the interface partition in the three-dimensional case.
@@
Roughly speaking, to recover the conductivity $\gamma$ on the current sub-region $L_{t+1}^{\mathcal{S}}\cup K_{t+1}^{\mathcal{S}}$, we select a set of boundary excitations that “focus” on the boundary of the sub-region $L_{t+1}^{\mathcal{S}}\cup K_{t+1}^{\mathcal{S}}$ and use its boundary measurements to finish the inductive step. The excitations are exactly in the kernel of the following map
@@
We can factorize the map $T^{(t)}$ into two parts: one is the map from the potential supported on the boundary $J_{t}^{\cal S}$ to the potential supported in the region $L_{t+1}$, and the other is the map from the potential supported on $L_{t+1}$ to the boundary currents on $\partial D\setminus J_{t}^{\cal S}$. We refer the reader to Fig. 2 for a schematic illustration in the three-dimensional case: the sought-for excitations are precisely supported on the blue nodes with the boundary current vanishing at yellow nodes. More precisely, we define the map $T_{1}^{(t)}:\mathbb{R}^{J_{t}^{\mathcal{S}}}\to\mathbb{R}^{L_{t+1}}$ by
@@
Then we denote the upper subgraph $G^{(t)}$ (i.e., above the interface $L_{t+1}$) by
$G^{(t)}=\bigl(E^{(t)},\,D\setminus L_{t+1}^{\mathcal{S}},\,L_{t+1}\cup(\partial D\setminus J_{t}^{\mathcal{S}})\bigr),$ (2.5)
where the set $E^{(t)}$ contains all edges in the set $E$ start or end at one node in $D\setminus L_{t}^{\mathcal{S}}$. In view of [19, Remark 3], the unique solvability of the discrete Dirichlet problem on the subgraph $G^{(t)}$ holds, and we denote by $S_{\gamma}^{(t)}$ the solution operator on the subgraph $G^{(t)}$. Hence the DtN map $\Lambda^{(t)}$ on the subgraph $G^{(t)}$ is well defined. We can now describe the map $T_{2}^{(t)}:\mathbb{R}^{L_{t+1}}\to\mathbb{R}^{\partial D\setminus J_{t}^{\mathcal{S}}}$, from the potential at $L_{t+1}$ to the current at $\partial D\setminus J_{t}^{\cal S}$, by a submatrix of $\Lambda^{(t)}$:
$T_{2}^{(t)}:=\Lambda^{(t)}(\partial D\setminus J_{t}^{\mathcal{S}};L_{t+1}).$ (2.6)
The next result rigorously establishes the factorization.
@@
For any $t$ in the range $d-1\leq t\leq dn-1$, the operator $T^{(t)}$ can be factorized as $T^{(t)}=T_{2}^{(t)}\circ T_{1}^{(t)}.$
@@
Fix any arbitrary $\varphi\in\mathbb{R}^{J_{t}^{\mathcal{S}}}$, and consider the zero extension of $\varphi$ to $\partial D$, still denoted by $\varphi$. The corresponding interior potential $\mathbf{u}$ is given by $S_{\gamma}\varphi$. By its definition, $\mathbf{u}$ is $\gamma$-harmonic on $G$, and also within the subgraph $G^{(t)}$. By the definition of the operator $T_{1}^{(t)}$, we have
@@
Since the discrete Dirichlet problem on $G^{(t)}$ has a unique solution, the restriction of $\mathbf{u}$ to the vertices of $G^{(t)}$ (i.e., $\overline{D}\setminus J_{t}^{\mathcal{S}})$ can be represented by
$\mathbf{u}|_{\overline{D}\setminus J_{t}^{\mathcal{S}}}=S_{\gamma}^{(t)}\bigl(T_{1}^{(t)}\varphi\bigr).$
Now we compute the boundary current and arrive at
@@
The factorization in Proposition 2.1 leads directly to the kernel inclusion relation:
@@
The identity $\ker T_{1}^{(t)}=\ker T^{(t)}$ then follows from the injectivity of $T_{2}^{(t)}$, which will be proved in Section 3.
If the boundary potential $\varphi$ is chosen from $\ker T_{1}^{(t)}$, the interior potentials are localized within the set $J_{t}^{\mathcal{S}}\cup L_{t}^{\mathcal{S}}$, since the Dirichlet boundary condition of the upper graph $G^{(t)}$ vanishes identically by the definition of the map $T_{1}^{(t)}$. We now define the solution spaces
@@
where $S_{\gamma}$ is the solution operator defined in (1.2). Note that, when recovering the conductivity on the edge set $E(L_{t}\cup K_{t},L_{t+1}\cup K_{t+1})$, only the boundary potentials in the quotient space
@@
provide active constraints, since for any $\varphi\in\ker T_{1}^{(t-1)}$ no current flows through $E(L_{t}\cup K_{t},L_{t+1}\cup K_{t+1})$. The quotient makes sense in view of the following inclusion result in the sense $\mathbb{R}^{J_{t-1}^{\mathcal{S}}}\subseteq\mathbb{R}^{J_{t}^{\mathcal{S}}}$.
@@
For any $\varphi\in\ker T_{1}^{(t-1)}$, let $\mathbf{u}:=S_{\gamma}\varphi$. By the construction,
$\mathbf{u}=0\quad\text{in }D\setminus L_{t-1}^{\mathcal{S}}\supseteq L_{t+1},$
@@
The monotonicity of $\mathcal{U}^{(t)}$ follows directly from Proposition 2.2: